OpenStudy (anonymous):

I'm doing an integrating by parts problem. u=cos(ln8x). What would du be?

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

\[\frac{ -\sin(\ln8x) }{ x }\]?

OpenStudy (anonymous):

\[-\sin\left(\ln 8x\right)\cdot \frac{1}{x} \] that's what i got

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."